Solve |5+x| > 7-3x.
Solve |x-2| - 3 ≤ 1.

Answer by ikleyn(52777)

.


I will solve the second equation, only.


The starting equation is

    | x - 2 | - 3 <= 1.


It is equivalent to

    | x - 2 | <= 1 + 3 = 4.


The last equation describes those values / (points in the number line) that

are remoted not far then 4 units from 2.


So the answer is  2 - 4 <= x <= 2 + 4,  or  -2 <= x <= 6.


In the interval form, the solution set is this segment on the number line 

    [-2,6].      ANSWER

Solved.